OpenStudy (anonymous):

Select the set of equations that represents the following situation: The sum of two numbers is 36, and their difference is nine more than the smaller number. What are the two numbers? 2y = 36; 36 = 9 – x x + y = 36; x – y = x – 9 x + y = 36; y – x = x + 9 2x = 36; 36 = x + 9

hero (hero):

Has the be the third option

OpenStudy (anonymous):

this one x + y = 36; y – x = x + 9

hero (hero):

Yes....because in y - x, x is portrayed as as the smaller number.

OpenStudy (anonymous):

k